SV Notes for MJD 56248 (starting evening of November 16)


Observer 1 Observer 2 Run Manager SISPI Expert on Call Eyeball Squad QR
Elliot Kent Gerdes Elliot Gruen/March Ogando

DECam Configuration and Setup

SISPI Configuration File sv.ini
Proposal ID 2012B-0003
Image Processing (Calibration) NCSA
Image Processing (First Cut) NCSA


Twilight ends 21:49, middle of night to hand over to community SV at 01:27

Moonset at 23:33 14% illuminated, 114 degrees away from our Mini-Survey field on Stripe 82. Still pretty dark!

Weather Conditions

Observing Plan and Goals for the Night

  1. Standard bias and dome flat sequence during afternoon
  2. Please try to get dome open for temperature equilibration at earliest appropriate time, be sure all cooling/ventilation systems are on.
Monitor Conditions:
  1. Check on DIMM status, running if at all possible. [Jiangang: I noticed the DIMM seeing changes over the past 3 days, but they change once and fixed there. See my attached image]
  2. Make sure RASICAM is online if weather conditions permit.
  3. A few times per night ask the operator to contact the other telescopes on Tololo to get their seeing value. Record those in the eLog.

Expectation for tonight is clear weather, decent seeing, tracking with modest ellipticity (<~0.1), poorer tracking N of zenith?

Note: SN observations are particular about the pointing, but it is ok to accept whatever pointing the telescope produces when doing standards and minisurvey.

Prioritization for tonight's observations:
  1. Photometric standards sequence as soon as dark enough (ugrizY in 3 fields at different airmasses, preferably one at X < 1.25, one at X = 1.25-1.65 and one at X > 1.65) and again at handover to Dara. Conditions: clear or thin clouds, any seeing.
  2. SPT-CL J2332-5358. Full-depth grizY observations. Conditions required: Clear or thin clouds, <1.6" seeing. Quit this field as soon as Minisurvey is observable unless images <1.2" are being obtained, since the whole point is to reach full depth on a field with better IQ than we are getting in North._
  3. Observe MiniSurvey under OBSTAC control, should be nearly completing gr exposures in MS1b as the MS1a izY exposures are still marked as "do not observe." OK to move on to MS1a/MS1b izY when the gr are completed. Conditions required: clear or thin clouds, seeing <1.6".
  4. DECal sweeps Conditions required: Any.

Notes from the 4PM Meeting and Additional Comments

Notes from Observers

See Night summary

Due to network problems last night QR was not running for a while and had to be restarted by Ann Elliot and Dara Norman.

TASCA b-filter animated GIF for this night (time stamp in UT):
TASCA b-filter animated GIF for this night

Notes on Analyses of Night's Data: